A car and driver, with a total mass of 1350kg, moving at 30m/s on a horizontal road when the driver sees roadworks ahead. She brakes, decelerating with a constant force of 600N until arriving at the roadworks. Find the time elapsed before arriving at the roadworks.


1
Expert's answer
2022-06-09T18:22:18-0400

"\vec F = m \vec a"

"\vec a =\frac{\vec F} {m} = \frac{600} {1350}\approx 0.44"

"\text{the car slows down consequently}"

"v_1 = v_0 -at"

"v_1 = 0"

"t = \frac{v_0}{a}= \frac{30}{0.44}\approx68.18 s"


"\text{Answer: } 68.18 s"